# Search Wisely: Mitigating Sub-optimal Agentic Searches By Reducing Uncertainty

### v1

@@ Abstract
Agentic Retrieval-Augmented Generation (RAG) systems enhance Large Language Models (LLMs) by enabling dynamic, multi-step reasoning and information retrieval. However, these systems often exhibit sub-optimal search behaviors like over-search (retrieving redundant information) and under-search (failing to retrieve necessary information), which hinder efficiency and reliability. This work formally defines and quantifies these behaviors, revealing their prevalence across multiple QA datasets and agentic RAG systems (e.g., one model could have avoided searching in 27.7% of its search steps). Furthermore, we demonstrate a crucial link between these inefficiencies and the models’ uncertainty regarding their own knowledge boundaries, where response accuracy correlates with model’s uncertainty in its search decisions. To address this, we propose $\beta$-GRPO, a reinforcement learning-based training method that incorporates confidence threshold to reward high-certainty search decisions. Experiments on seven QA benchmarks show that $\beta$-GRPO enable a 3B model with better agentic RAG ability, outperforming other strong baselines with a 4% higher average exact match score [^1] We will release all our codes and data upon acceptance..
@@ 2.1 Step-wise Analysis
To directly measure whether a search step was truly necessary, we separate all outputs into individual steps and identify if each of them matches with the definition of over-search and under-search. For over-search rate measurement, we prompted the model to answer sub-queries from all the steps with search behavior using only their internal knowledge and the preceding context. For under-search rate measurement, we examine steps without searching and evaluate the correctness of the generated information. A detailed explanation of the analysis pipeline is provided in Appendix A.2.
@@ 2.2 Sub-optimal Search & Knowledge Boundary
The observed tendencies towards over-search and under-search, combined with our definition, suggest a core deficiency in how agentic RAG models perceive knowledge boundaries—the limits of what they know versus what they need to find out. To illustrate the link between better knowledge boundary awareness and improved outcomes, we analyze the performance of 4 Qwen2.5-3B based Search-R1 models (including PPO and GRPO trained, Base and Instruct variants). We generate 5 candidate responses for each question and group these responses based on each output’s minimum probabilities within all the search query tokens as the indication of certainty on knowledge boundary.
@@ 3 Approach
$\beta$-GRPO Motivated by the observation that rollouts with low-confidence search calls are more likely to be incorrect, we incorporate model confidence into the RL reward process. Specifically, for each rollout containing search calls (enclosed within <search></search> tags), we extract the probabilities of the search tokens including the tags and use the minimum probability among them as a measure of the model confidence for the search calls within a rollout Jiang et al. (2023). We then set a confidence threshold $\beta$: only rollouts with the confidence of search calls (if exist) above $\beta$ and correct answers receive a reward of 1, otherwise 0.
@@ 4 Experiments
Search-R1-GRPO Search-R1-$\beta$-GRPO (ours)
Question: In what Country is Sul America Esporte Clube in? Generation: <think> The question asks for the country hosting Sul America Esporte Clube, a team in the brazilian C1 and C2 league, thus the direct reasoning is likely to be Brazilian football team residing in Brazil. Question: In what Country is Sul America Esporte Clube in? Generation: <search> in what Country is Sul America Esporte Clube in </search> <information> Doc 1: … Doc 2: … Doc 3: … </information> <answer> Brazil </answer> (Confidence: 0.99)
Table 3: Case Study. The retrieved documents are omitted for clarity.
@@ 5 Analysis
Ablation on $\beta$ Case Study Following Jiang et al. (2023), we experiment with three confidence threshold values: 0.2, 0.4, and 0.6. The average EM scores are 0.341, 0.344 and 0.336 with a threshold of 0.4 yields the best result. Moreover, we find 115 test cases from the multi-hop QA datasets where Search-R1-$\beta$-GRPO produces a correct answer with higher confidence, while Search-R1-GRPO gives an incorrect answer. These cases clearly benefit from the increased model confidence enabled by the proposed $\beta$-GRPO. An example is shown in Table 3: Search-R1-GRPO lacks confidence and fails to provide a definite answer, whereas Search-R1-$\beta$-GRPO generates a confident search query and produces the correct answer.
@@ 6 Conclusion
In this work, we formally define and quantify sub-optimal search behaviors, over-search and under-search, in agentic RAG systems, revealing their prevalence and impact. By introducing $\beta$-GRPO, a confidence-aware policy gradient method, we enable a 3B model with better agentic RAG ability than strong baselines.
7 Limitations
@@ Appendix A Appendix
A.1 Formal Definition of Under-search & Over-search
A.2 Detailed Step-wise Analysis Procedure
@@
    Step Extraction: We parse the agent’s interaction log following the definition in Appendix A.1. Each distinct thinking process is a decision point and considered a step, typically delineated by <step> and </step> tags (or a similar structured logging format). A "search step" is identified as any step where all three relevant operations—think (the model’s reasoning), search (the search query issued), context (the information retrieved). A "non-search step" typically only consists of thinking. In this work specifically, the Step Extraction is done by prompting QwQ-32B Qwen et al. (2025) as we discover that reasoning LLM typically perform better on such task.
@@
A.3 Search Frequency vs. Optimal Hops

### v2

@@ Abstract
Agentic Retrieval-Augmented Generation (RAG) systems enhance Large Language Models (LLMs) by enabling dynamic, multi-step reasoning and information retrieval. However, these systems often exhibit sub-optimal search behaviors like over-search (retrieving redundant information) and under-search (failing to initiate retrieval for necessary information), which hinder efficiency and reliability. This work formally defines and quantifies these behaviors, revealing their prevalence across multiple QA datasets and agentic RAG systems (e.g., one model could have avoided searching in 27.7% of its search steps). Furthermore, we demonstrate a crucial link between these inefficiencies and the models’ uncertainty regarding their own knowledge boundaries, where response accuracy correlates with model’s uncertainty or confidence in its search decisions. To address this, we propose $\beta$-GRPO, a reinforcement learning-based training method that incorporates confidence threshold to reward high-certainty search decisions. Experiments on seven QA benchmarks show that $\beta$-GRPO enable a 3B model with better agentic RAG ability, outperforming other strong baselines with a 4% higher average exact match score, with lower over-search and under-search rate [^1] We have release our code for training at https://github.com/mianzhang/Search-R1..
@@ 2 Identifying Sub-optimal Search
2.1 Formal Definition of Under-search & Over-search
@@ 2.2 Step-wise Analysis
To directly measure whether a search step was truly necessary, we separate all outputs into individual steps and identify if each of them matches with the definition of over-search and under-search as described in Section 2.1. For over-search rate measurement, we prompted the model to answer sub-queries from all the steps with search behavior using only their internal knowledge and the preceding context. For under-search, we examine steps without searching and evaluate the correctness of the generated information. A detailed explanation of the analysis pipeline is provided in Appendix A.1 with a flow chart in Figure 1.
@@
Figure 1: Flowchart of analysis pipeline for over-search and under-search.
@@ 2.3 Sub-optimal Search & Knowledge Boundary
The observed tendencies towards over-search and under-search, combined with our definition, suggest a core deficiency in how agentic RAG models perceive knowledge boundaries—the limits of what they know versus what they need to find out. To illustrate the link between better knowledge boundary awareness and improved outcomes, we analyze the performance of 4 Qwen2.5-3B based Search-R1 models (including PPO and GRPO trained, Base and Instruct variants). We generate 5 candidate responses for each question and group these responses based on each output’s minimum probabilities within all the search query tokens in a trajectory as the indication of certainty on knowledge boundary.
@@ 3 Approach
$\beta$-GRPO Motivated by the observation that rollouts with low-confidence search calls are more likely to be incorrect, we incorporate model confidence into the RL reward process. Specifically, for each rollout containing search calls (enclosed within <search></search> tags), we extract the probabilities of the search tokens including the tags and use the minimum probability among them as a measure of the model confidence for the search calls within a rollout Jiang et al. (2023). We then set a confidence threshold $\beta$: only rollouts with the confidence of search calls (if exist) above $\beta$ and correct answers receive a reward of 1, otherwise 0. Formally, for a given reasoning trajectory $T={s_{1},s_{2},...,s_{N}}$ as described in Section 2.1, with predicted final answer $a_{f}$, ground-truth final answer $a^{*}_{f}$, and confidence $C(T)$ the reward $R(T)$ is calculated as Equation 1,
$R(T)=\begin{cases}1,&\text{if $a_{f}=a^{*}_{f}\wedge C(T)>\beta$}\\ 0,&\text{otherwise}\end{cases}$ (1)
where $C(T)$ is the confidence of trajectory $T$ calculated as Equation 2,
$C(T)=\min_{s_{t}^{R}\in T,w\in q_{t}}P(w)$ (2)
where $w$ is the token that makes up the search query $q_{t}$ for a given retrieval step and $P(w)$ is the probability assigned to $w$.
@@ 4 Experiments
Baseline Model (Search-R1-GRPO) Our Model (Search-R1-$\beta$-GRPO)
Case 1: Mitigating Over-search for common knowledge questions
Question: Who was the first president of the United States? Question: Who was the first president of the United States?
Generation: <think> The user is asking a question about a historical figure. To ensure accuracy, I will perform a search. </think> <search> first president of the United States </search> Generation: <think> The user is asking for the first president of the United States. This is a well-established fact stored in my parametric knowledge. I am highly confident in the answer and do not need to perform an external search. </think>
<information>Doc 1: … Doc 2: … </information> <answer> George Washington </answer>
<answer> George Washington </answer>
Correctness: Correct Correctness: Correct
Case 2: Mitigating Under-search for knowledge-intensive questions
Question: In what Country is Sul America Esporte Clube in? Question: In what Country is Sul America Esporte Clube in?
Generation: <think> The question asks for the country of Sul America Esporte Clube. The name "Sul America" sounds Portuguese or Spanish, suggesting a location in South America. It could be in Argentina or Colombia. I will make a guess based on common football leagues. </think> Generation: <think> The user is asking for the location of a specific sports club, "Sul America Esporte Clube". This is a niche entity, and my internal knowledge is likely insufficient. To provide an accurate answer, I must perform a search. </think> <search> country of Sul America Esporte Clube </search>
<answer> Argentina </answer> <information> Doc 1: … Doc 2: … </information>
<answer> Brazil </answer>
(Confidence: 0.99)
Correctness: Incorrect Correctness: Correct
Table 3: Case Study comparing search behaviors. The top pair demonstrates how our model avoids an unnecessary search (over-search) for a common fact. The bottom pair shows our model correcting a failure to search (under-search), preventing hallucination for a niche query. Retrieved documents are omitted or summarized for clarity.
@@ 5 Analysis
Ablation on $\beta$ & Case Study Following Jiang et al. (2023), we experiment with three confidence threshold values: 0.2, 0.4, and 0.6. The average EM scores are 0.341, 0.344 and 0.336 with a threshold of 0.4 yields the best result. Moreover, we find 115 test cases from the multi-hop QA datasets where Search-R1-$\beta$-GRPO produces a correct answer with higher confidence, while Search-R1-GRPO gives an incorrect answer. These cases clearly benefit from the increased model confidence enabled by the proposed $\beta$-GRPO.
@@
Case Study Our case study in Table 3 highlights the model’s improved search decisions. For a simple question ("Who was the first president?"), the baseline model performs an unnecessary search, whereas our model confidently answers from its internal knowledge. Conversely, when faced with an obscure query ("In what Country is Sul America Esporte Clube in?"), the baseline hallucinates an incorrect answer. Our model correctly identifies this knowledge gap, initiates a search, and provides the accurate answer.
@@ 6 Conclusion
In this work, we formally define sub-optimal search behaviors, over-search and under-search, in agentic RAG systems. Our analysis showed these behaviors are widespread; for instance, one model could have avoided searching in 27.7% of its search steps, while another exhibited error rates as high as 63% in non-search steps, indicating significant under-searching. We established a link between these inefficiencies and a model’s uncertainty about its knowledge boundaries, finding that higher confidence in search decisions correlates with better accuracy. By introducing $\beta$-GRPO, a confidence-aware policy gradient method, we enable a 3B model with better agentic RAG ability than strong baselines. This approach, which rewards only high-certainty search decisions that lead to correct answers, resulted in a 4% higher average exact match score and notable reductions in both over-search and under-search rates. Future work should explore more sophisticated and fine-grained reward design on trajectory, with experiments on larger size models.
Limitations
@@
Acknowledgements
We thank the anonymous reviewers for their thoughtful comments and suggestions. We are also grateful to our colleagues and collaborators for their insightful discussions and support throughout this project.
@@ Appendix A Appendix
A.1 Detailed Step-wise Analysis Procedure
@@
    Step Extraction: We parse the agent’s interaction log following the definition in Appendix 2.1. Each distinct thinking process is a decision point and considered a step, typically delineated by <step> and </step> tags (or a similar structured logging format). A "search step" is identified as any step where all three relevant operations—think (the model’s reasoning), search (the search query issued), context (the information retrieved). A "non-search step" typically only consists of thinking. In this work specifically, the Step Extraction is done by prompting QwQ-32B Qwen et al. (2025) as we discover that reasoning LLM typically perform better on such task. An example of instruction for step extraction is provided in Appendix A.4.
@@
A.2 Search Frequency vs. Optimal Hops
@@ A.3 Instruction for Model Input
A.4 Instruction for Step Extraction
Objective: Your task is to parse a complete interaction log from a reasoning agent. Your goal is to segment the log into a chronological sequence of steps and structure the output for each step into a consistent JSON format.
Core Methodology: A reasoning trajectory is a sequence of steps, $s_{1},s_{2},\dots,s_{N}$. Each step, $s_{t}$, involves a reasoning component, $r_{t}$. The sub-answer, $a_{t}$, which is the conclusion of step $s_{t}$, is reflected in the reasoning component of the next step, $r_{t+1}$. You must follow this look-ahead method to determine the conclusion for each step.
Key Definitions:
Search Step:
A step where the agent uses a search tool. It must include a reasoning block (<think>), a search query (<search>), and retrieved context (<information>).
Non-Search Step:
A step where the agent relies only on its internal knowledge and prior context. It typically only includes a reasoning block (<think>).
Conclusion:
The sub-answer or piece of information the agent generates or confirms at the end of a step ($s_{t}$), which contributes to the final answer.
Instructions:
  1. 1.
    Parse the entire interaction log into a chronological sequence of steps. A new step begins with each distinct reasoning block (e.g., content within <think> tags).
  2. 2.
    For each step ($s_{t}$), extract the following components:
    • reasoning: The content from the <think> block of the current step.
    • query: The content from the <search> block of the current step. If not present, use null.
    • information: The content from the <information> block of the current step. If not present, use null.
    • conclusion: The specific sub-answer ($a_{t}$) produced as a result of the current step’s actions. Crucially, you must identify this conclusion by analyzing the reasoning block of the following step ($r_{t+1}$), where it is first used or stated. For the final step in the trajectory, the conclusion is the final answer itself.
  3. 3.
    For each step, construct a JSON object containing the extracted components.
  4. 4.
    Present the final output as a sequence of <step> blocks, with each block containing the JSON object for that step.
Required Output Format: Your entire output must be a sequence of <step> blocks.
Example for a Search Step:
1<step>
2{
3 "reasoning": "The user is asking for the capital of France. I should search for this information to be certain.",
4 "query": "capital of France",
5 "information": "Paris is the capital and most populous city of France...",
6 "conclusion": "The capital of France is Paris."
7}
8</step>
Example for a Non-Search Step:
1<step>
2{
3 "reasoning": "Now that I know the capital is Paris, I can formulate the final answer.",
4 "query": null,
5 "information": null,
6 "conclusion": "The final answer is Paris."
7}
8</step>